OpenStudy (mathmath333):

is \((a-b)!=a!-b!\) ? ,\( a\geq 1,b\geq 1\)

hartnn (hartnn):

! is factorial ?

OpenStudy (anonymous):

@hartnn yes

OpenStudy (anonymous):

@hartn can you help me with a question after you are done

hartnn (hartnn):

if its factorial, then that statement is not true counterexample : let a =4, b =3 (4-3)! =1 4! -3! is NOT 1

hartnn (hartnn):

also if b>a then that statement is definitely false, as factorial of negative numbers is not defined

OpenStudy (mathmath333):

so is it true for \(\Large a>b>1\)

hartnn (hartnn):

nopes, even of a>b>1 say a = 5 , b =3 (5-3)! = 2! =2 5! -3! is not 2 infact, (a-b)! < a!-b!

hartnn (hartnn):

or i should say \((a-b)! \le a!-b! \) except for a,b =1

OpenStudy (mathmath333):

ok but somehow if i want to convert \(\Large (a-b)!\implies a!-b!\) then is their a way to do it

hartnn (hartnn):

\((a-b)!\) cannot be simplified

ganeshie8 (ganeshie8):

are you trying to telescope something ?

OpenStudy (mathmath333):

example\(\Large (5-3)!=5!-3!+C\)

ganeshie8 (ganeshie8):

whats the original complete question ?

OpenStudy (mathmath333):

well it was CAT paper today , there was a question like that i cant remeber

ganeshie8 (ganeshie8):

oh you gave the exam today ? how was it ?

OpenStudy (mathmath333):

it was easy but tricky ,math section was easy , and time consuming , even in english section they ask math question

ganeshie8 (ganeshie8):

lol yeah data terpretation takes long time ! good to knw you did well :)

OpenStudy (mathmath333):

yes i learned that a speed basic tricks and common sense needed, i think any could have solved it but it is very time consuming ,50 questions for 1 hr and 25 min

OpenStudy (mathmath333):

40% were from number theory

hartnn (hartnn):

let us know if you recall the exact question :)